The probability of event A=0.15 and event B=0.25, what is the probability of A and B both occuring?

a. 0.15*0.25=0.0375

b. 0.15+0.25 = 0.40

c. Can't tell because we don't know if A and B are independent events
Answer #1

The correct option is A

Explanation:

Given probability of event A=0.15 and event B=0.25,

then the probabilty of P(A n B) = P(A)*P(B) = 0.15 * 0.25 = 0.0375